OpenStudy (anonymous):

Find the slope of the line passing through each pair of points. (-3, 4) and (6, 1)

OpenStudy (shamil98):

Use the slope formula. \[\huge \frac{ y_2 - y_1 }{ x_2 - x_1 }\]

OpenStudy (anonymous):

C o m e h e l p m e!

OpenStudy (anonymous):

4 - 1 _______ -3 - 6 ^ Like this...?

OpenStudy (shamil98):

No, you got the coordinates mixed up. try again

OpenStudy (shamil98):

it should be the second y - coordinated minus the first one

OpenStudy (shamil98):

amd the same goes for the x coordinates

OpenStudy (anonymous):

1 - 4 _______ 6 - -3

OpenStudy (shamil98):

Yep, now simplify it

OpenStudy (anonymous):

-3 ___ 9

OpenStudy (shamil98):

correct, and there you go -3/9 is your slope

OpenStudy (anonymous):

it doesn't really matter which coordinate your subtracting from what btw

OpenStudy (shamil98):

yes it does.

OpenStudy (anonymous):

They gave me an answer key (without the explanation) and they said the answer is -1/3... Is that -3/9 simplified?

OpenStudy (anonymous):

it does not matter.

OpenStudy (shamil98):

yes that is

OpenStudy (shamil98):

Why do you think there is formula for it?

OpenStudy (shamil98):

You're mislead..

OpenStudy (anonymous):

(1)-(4)/(6)-(-3) = -3/9 (4)-(1)/(-3)-(6) = -3/9 clearly, i'm misled.

OpenStudy (shamil98):

You are mislead, your way won't work for every single pair of coordinates.

OpenStudy (shamil98):

Sure you get the same answer this time but in other cases you wont

OpenStudy (anonymous):

show me a case where it wont.